ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 57,58 Shubladze Mamuka, Georgian Technical University, Tbilisi, Georgia The topology of hypersurface singularities with codimension one let {f = 0} be a hypersurface in C^{n+1} with 1-codimensional singular set Σ wich has a transversal singularity of type A_{k−1}. For this class of singularities we investigate criterion of finite multiplicity, constract special deformations which preserves topology structure of Milnor fibre, calculate the number of Morse points in terms of the corespoding local rings, and describe homotopy type of the Milnor fibre.